Please Solve. –3(x + 1) = 2(x – 8) + 3
–3(x + 1) = 2(x – 8) + 3 [Use the Order of Operations to solve (PEMDAS)]
-3x-3=2x-16+3 [Combine like-terms if possible]
-3x-3=2x-13 [Isolate all of the terms with variables]
-3x-2x=-13+3 [Combine like-terms]
-5x=-10 [Solve for x]
x=-10/-5
x=2

Found 2 solutions by tutorcecilia, TakeATuition.com:
Answer by tutorcecilia(2152)

–3(x + 1) = 2(x – 8) + 3 [Use the Order of Operations to solve (PEMDAS)]
-3x-3=2x-16+3 [Combine like-terms if possible]
-3x-3=2x-13 [Isolate all of the terms with variables]
-3x-2x=-13+3 [Combine like-terms]
-5x=-10 [Solve for x]
x=-10/-5
x=2
.
.
Check by plugging (x=2) back into the original equation.

Answer by TakeATuition.com(57)